OpenStudy (mackenzie2013):

Terrance has $100 that he can invest in a savings account. His bank is willing to pay interest to his account at a rate of 0.04. How long will Terrance need to leave the money in the account if he wants to earn $100 in interest, doubling his money? Use the formula . T= I/PR 25 years 10 years 40 years 5 years

OpenStudy (anonymous):

so first you need to put the numbers in to match the formula ya know what i mean?

OpenStudy (mackenzie2013):

Yes but im confused

OpenStudy (anonymous):

what makes you confused?

OpenStudy (mackenzie2013):

I know the rate is 0.04 I is the amount of investment which is 100$ Thats really all I understand...probably wrong to lol

OpenStudy (mackenzie2013):

@JesseJ0906

OpenStudy (anonymous):

ok does T mean time?

OpenStudy (mackenzie2013):

Yes.

OpenStudy (anonymous):

ok what is P? i can't remember

OpenStudy (mackenzie2013):

P represents the principle or amount invested

OpenStudy (anonymous):

ok so it'd be T=100/100*.04

OpenStudy (anonymous):

jeez i just realized i say ok A LOT

OpenStudy (mackenzie2013):

So it would be 0.04?

OpenStudy (anonymous):

no hah it is 25 because .04*100=4 amd 100/4=25

OpenStudy (mackenzie2013):

OK so you take .04 and times it by 100 ?

OpenStudy (anonymous):

yep i think i just forgot to put parenthesis somewhere i tend to do that

OpenStudy (mackenzie2013):

and then you take 4 and divide it by 100?

OpenStudy (anonymous):

yep

OpenStudy (mackenzie2013):

thanks !

OpenStudy (anonymous):

np :)